A car is moving towards south with a speed of 20m/s. A motor cyclist is moving towards east with a speed of 15m/s. At a certain instant, the mortcyclist is due south of the car and is at a distance of 50m form the car. Find the shortest distance between the motorcyclist and the car. Also find the time after which they are closest to each other.

Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

 

Dear Prateek, 

let t be the time from the instant when the bike is due south. distance squared = (50-20t)2 +(15t)2= 625t2-2000t +2500

='s'. ds/dt=0, 1250t-2000=0, tmin=1.6 seconds, d2s/dt2=1250t=1250*1.6>0, minimum distance=√s at t=1.6 =√900=30m

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar.. 

Askiitians Expert

Chetan Mandayam Nayakar – IIT Madras